OpenStudy (anonymous):

Solve: 3^(4x-7) = 4^(2x+3)

OpenStudy (anonymous):

start with \[(4x+7)\ln(3)=(2x+3)\ln(4)\] and then do a bunch of algebra to solve for x

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

yw that was the easy part, algebra is the hard part